# PerceiverS: A Multi-Scale Perceiver with Effective Segmentation for Long-Term Expressive Symbolic Music Generation

### v2

@@ Abstract
AI-based music generation has progressed significantly in recent years. However, creating symbolic music that is both long-structured and expressive remains a considerable challenge. In this paper, we propose PerceiverS (Segmentation and Scale), a novel architecture designed to address this issue by leveraging both Effective Segmentation and Multi-Scale attention mechanisms. Our approach enhances symbolic music generation by simultaneously learning long-term structural dependencies and short-term expressive details. By combining cross-attention and self-attention in a Multi-Scale setting, PerceiverS captures long-range musical structure while preserving musical diversity. The proposed model has been evaluated using the Maestro dataset and has demonstrated improvements in generating music of conventional length with expressive nuances. The project demos and the generated music samples can be accessed through the link: https://perceivers.github.io.
Index Terms:
Symbolic Music Generation, Long-Term Structure, Transformer Models, Cross-Attention, Self-Attention, PerceiverS, Effective Segmentation, Multi-Scale AttentionI Introduction
@@
Recent advancements in AI-based music generation, especially in audio generation models such as AudioLDM [1], MusicGen [2], and Jen-1 [3], have demonstrated significant progress capable of generating highly natural-sounding music. Compared to audio generation, symbolic music offers a further level of abstraction, making it easier for machine learning models to capture deeper musical characteristics and understanding. Symbolic music plays a crucial role in the field of music generation, particularly due to its editable nature. This allows for operations such as cutting and rearranging different sections or substituting instrument timbres, enabling human involvement in high-quality music production during the post-processing stage.
@@
However, symbolic music generation faces two key challenges. First, the most expressive datasets, recorded from live performances and recording studios, are seldom used compared to manually created MIDI-file datasets. The primary reason is that they lack detailed annotations, making it harder for models to learn their complex structures. Furthermore, due to computational limitations, previous models cannot fully capture the context of an entire piece of music. Techniques, such as chunking and quantization, are often employed to reduce computational complexity, leading to the loss of crucial musical details and making it difficult for models to grasp the full structure of a composition. Second, other approaches that use abstract structural representations as conditions for music generation tasks have enabled the generation of structured music. However, such methods rely heavily on handcrafted feature engineering. Our objective is to design and develop a model that is capable of learning the long-range dependencies in music without relying on explicit structural annotations.
@@
The emergence of Transformer Attention technologies, such as Perceiver AR [4], has made it possible to access much longer contextual dependencies. It allows for the simultaneous learning of musical structure and the generation of expressive performances. Perceiver AR has demonstrated the ability to attend to a context length of up to 32,768 tokens using the Maestro dataset [5], where the query in cross-attention attends to significantly longer key/value pairs [4]. However, the causal mask, when applied with the default input sequence segmentation, does not fully conceal tokens that should not be visible during auto-regressive training and generation, which ultimately degrades the quality of the generated music. Additionally, when using ultra-long context as a single condition, the model tends to generate identical or similar repetitive segments as the sequence length increases due to issues with high similarity in the context of neighboring tokens, leading to a high token autocorrelation [6] tendency.
@@
In this paper, we propose PerceiverS, a novel model that addresses the causal masking issue by incorporating Effective Segmentation. Additionally, PerceiverS employs Multi-Scale attention to mitigate the high token autocorrelation problem that arises from relying solely on long-range dependencies. Specifically, by adjusting the input sequence segmentation to start from the head segment with an effective causal mask and aggressively increasing the segment length up to the maximum input sequence length, we resolve the learning limitations caused by the causal mask in Perceiver AR [4]. Additionally, by incorporating Multi-Scale masks across multiple layers of cross-attention, the model considers both ultra-long and short-range attention simultaneously. This approach addresses the limitation in Perceiver AR, which focuses solely on long-range attention [4]. Different from Perceiver AR, our approach enhances symbolic music generation by effectively capturing long-term structural dependencies while simultaneously focusing on short-term variations. Through improved segmentation and multi-scale attention mechanisms, PerceiverS generates coherent, diverse music without relying on extensive structural annotations. Extensive experiments have been conducted to evaluate the performance of the proposed PerceiverS. The experimental results demonstrate an average 40% improvement in Overlap Area when measured against the original training dataset, highlighting a substantial advantage of our approach over Perceiver AR [4] in generating high-quality symbolic music.
@@ II-A1 Dataset Selection
A key factor in generating production-quality music lies in the selection of datasets. The Maestro dataset [5], consisting of real human performances, offers dynamic and expressive recording music. Datasets recorded from live performances are rare, but there are quite a few Automatic Music Transcription (AMT) datasets, including GiantMIDI [8], ATEPP [9], PiJAMA [10], and others. Advanced models have been developed, from Hawthorne [11] to Kong [8], that convert audio into MIDI files. These advances have made it possible to use a vast amount of recorded audio music to train models, as the development of symbolic music models has long been constrained by the limited availability of datasets.
Manually created MIDI datasets, like LAKH [12], provide valuable human-annotated information, e.g., beats, bars, and phrases, which allows for flexible segmentation and richer feature extraction but lack expressive nuances found in live performances, such as dynamics, tempo variations, and subtle timing shifts. Other attempts, such as ASAP [13] with human-assisted beat correction, and advancements in automatic beat tracking, such as Beat This! [14], aim to bring annotation to live-recorded datasets, though accuracy limitations still present challenges.
@@ II-B1 Explicit Structural Features
The explicit use of structural features often relies on handcrafted feature engineering and external analyzing tools. A common method in these models is a waterfall-like approach. Typically, the process begins by generating a lead sheet and then using the lead sheet as a condition for the subsequent generation tasks.
@@ II-B2 Implicit Structural Features
II-C SOTA Solutions with Long-Term Dependency On Performance Datasets
@@
However, Perceiver AR leverages the last tokens as the Query with a limited causal mask, and training with teacher-forcing on long sequences led to lower quality generation. Furthermore, we observed that relying solely on long, especially ultra-long, context resulted in repetitive segments in the latter part of generated content.
Despite significant progress, existing research has yet to fully address the challenge of generating music with both long-term dependencies and expressive performance details, as most approaches either rely heavily on manual processes or face performance limitations due to restricted context window lengths. While Perceiver AR represents the state-of-the-art method for generating music with long-term dependencies, there remain areas that require further improvement.
@@ III-A2 Perceiver AR’s Causal Mask Issue
In Perceiver AR [4], the situation is different because the query token length is much smaller than the key and value lengths . Specifically, the causal mask only works on the final part of the context sequence, equivalent to the length of the query , but does not mask tokens that occur before that. This results in some tokens before the query length being visible during training, which is not an issue for generation except that the segment from to is not properly learned by the model.
@@
Note that this issue primarily affects unconditional generation, where the model generates music without any specific prompt or primer, making it more sensitive to inconsistencies between training and generation contexts. In conditional generation, where the model starts with an initial prompt or primer sequence, using a primer with a length close to the unmasked portion during training can help preserve generation quality and prevent degradation.
@@ III-B Ultra-Long Context in auto-regressive Generation
During generation, identical values are not produced at nearby steps to avoid training penalties. In fact, the similar context of neighboring tokens causes the generation process to produce identical or similar tokens at nearby steps, leading to a higher probability of generating repetitive short segments as the sequence grows longer.
@@ IV-A Improving the Model to Effectively Learn Ultra-Long Sequences
The upper part of Figure 1 demonstrates the original input sequence pre-processing of the baseline model. Specifically, the baseline model segments the input sequence using the maximum input length as the window size, leaving the beginning of the sequence uncovered by the causal mask. This results in these initial tokens not being progressively used as validation tokens in teacher forcing, preventing the model from learning to generate tokens at the start of the sequence that fall outside the mask.
@@ V-A Effective Segmentation
This approach ensures that the model learns the sequence structure progressively, closely resembling the auto-regressive generation process where tokens are predicted one at a time. As a result, the model is better aligned with the auto-regressive nature of music generation, reducing the peeking problem and improving generation quality. Thus, the model can smoothly learn the generation of sequences with an effective causal mask. The key advantages are summarized as follows.
@@ V-B Multi-Scale Cross-Attention Mechanism
First, we introduce the Attention mask without scale mask. In this case, all tokens are visible to the model. as defined in Equation (15):
@@ VI-A3 Hyperparameter Selection
The hidden dimension was set to 1,024 with 24 self-attention layers. Each attention layer had 16 heads, and the head dimension was 64. Adam optimization was used, with an initial learning rate set to 0.03125. Each generated sequence length was set to 8,192 tokens, resulting in approximately 2-10 minutes of music. The music generation in this setup was unconditional, meaning that no external conditions or prompts were used to guide the generation process. The resulting MIDI files were rendered into audio using the Vintage Piano sound from Logic Pro.
@@ VI-A4 Evaluation Metrics
To calculate these additional segment-based metrics, each evaluation piece was divided into 64 segments of equal duration. All of these metrics are calculated based on the inter-set distribution similarity between the generated dataset and the ground truth dataset, including two values: the KL Divergence (KLD) and Overlap Area (OA). When the OA value is larger and the KLD value is smaller, it indicates that the above metric distributions of the generated dataset and the ground truth dataset are closely aligned, suggesting that the model has generated music more closely resembling human-composed and performed pieces.

### v3

@@ Abstract
AI-based music generation has made significant progress in recent years. However, generating symbolic music that is both long-structured and expressive remains a significant challenge. In this paper, we propose PerceiverS (Segmentation and Scale), a novel architecture designed to address this issue by leveraging both Effective Segmentation and Multi-Scale attention mechanisms. Our approach enhances symbolic music generation by simultaneously learning long-term structural dependencies and short-term expressive details. By combining cross-attention and self-attention in a Multi-Scale setting, PerceiverS captures long-range musical structure while preserving performance nuances. The proposed model has been evaluated using the Maestro dataset and has demonstrated improvements in generating coherent and diverse music, characterized by both structural consistency and expressive variation. The project demos and the generated music samples can be accessed through the link: https://perceivers.github.io.
I Introduction
@@
Recent advancements in AI-based music generation, especially in audio generation models such as AudioLDM [1], MusicGen [2], and Jen-1 [3], have demonstrated significant progress capable of generating highly natural-sounding music. Compared to audio generation, symbolic music offers a higher level of abstraction, making it easier for machine learning models to capture and understand deeper musical characteristics. Symbolic music plays a crucial role in the field of music generation, particularly due to its editable nature. This allows for operations such as cutting and rearranging different sections or substituting instrument timbres, enabling human involvement in high-quality music production during the post-processing stage.
@@
However, symbolic music generation faces two key challenges. First, the most expressive datasets, recorded from live performances and recording studios, are seldom used compared to manually created MIDI-file datasets. The primary reason is that they lack detailed annotations, making it harder for models to learn their complex structures. Furthermore, due to computational limitations, these models cannot fully capture the context of an entire piece of music. Techniques, such as chunking and quantization, are often employed to reduce computational complexity, leading to the loss of crucial musical details and making it difficult for models to grasp the full structure of a composition. Second, other approaches that use abstract structural representations as conditions for music generation tasks have enabled the generation of structured music. However, such methods rely heavily on handcrafted feature engineering. Our objective is to design and develop a model that is capable of learning the long-range dependencies in music without relying on explicit structural annotations.
@@
The emergence of Transformer Attention technologies, such as Perceiver AR [4], has made it possible to access much longer contextual dependencies. It allows for the simultaneous learning of musical structure and the generation of expressive performances. Perceiver AR has demonstrated the ability to attend to a context length of up to 32,768 tokens using the Maestro dataset [5], where the query in cross-attention attends to significantly longer key/value pairs [4].
In the Transformer attention architecture, a causal mask is a key mechanism that prevents each token from accessing future tokens during training. This prevents information leakage and enforces autoregressive behavior. A detailed explanation of this mechanism is provided in Section III.
However, the causal mask, when applied with the default input sequence segmentation, does not fully conceal tokens that should not be visible during auto-regressive training and generation, which ultimately degrades the quality of the generated music. Additionally, when using ultra-long context as a condition, the model tends to generate identical or similar repetitive segments as the sequence length increases due to issues with high similarity in the context of neighboring tokens, leading to a high token autocorrelation [6] tendency.
@@
In this paper, we propose PerceiverS, a novel model that addresses the causal masking issue by incorporating Effective Segmentation. Additionally, PerceiverS employs Multi-Scale attention to mitigate the high token autocorrelation problem that arises from relying solely on long-range dependencies. Specifically, by adjusting the input sequence segmentation to start from the head segment with an effective causal mask and aggressively increasing the segment length up to the maximum input sequence length, we resolve the learning limitations caused by the causal mask in Perceiver AR [4]. Additionally, by incorporating Multi-Scale masks across multiple layers of cross-attention, the model considers both ultra-long and short-range attention simultaneously. This approach addresses the limitation in Perceiver AR, which focuses solely on long-range attention [4]. Different from Perceiver AR, our approach enhances symbolic music generation by effectively capturing both long-term structural dependencies and short-term expressive details. Through improved segmentation and multi-scale attention mechanisms, PerceiverS generates coherent, diverse music without relying on extensive structural annotations. Extensive experiments have been conducted to evaluate the performance of the proposed PerceiverS. The experimental results demonstrate an average 40% improvement in Overlap Area when measured against the original training dataset, highlighting a substantial advantage of our approach over Perceiver AR [4] in generating high-quality symbolic music.
@@ II-A1 Dataset Selection
A key factor in generating production-quality music lies in the selection of datasets. The Maestro dataset [5], consisting of real human performances, offers dynamic and expressive recording music. Datasets recorded from live performances are rare, but there are quite a few Automatic Music Transcription (ATM) datasets, including GiantMIDI [8], ATEPP [9], PiJAMA [10], and others. Advanced models have been developed, from Hawthorne [11] to Kong [8], that convert audio into MIDI files. These advances have made it possible to use a vast amount of recorded audio music to train models, as the development of symbolic music models has long been constrained by the limited availability of datasets.
However, existing publicly available performance datasets are limited to solo piano recordings. There are currently no open datasets featuring solo performances of other instruments or multi-instrument ensemble performances.
Manually created MIDI datasets, like LAKH [12], provide valuable human-annotated information, e.g., beats, bars, and phrases, which allows for flexible segmentation and richer feature extraction but lack expressive nuances found in live performances, such as dynamics, tempo variations, and subtle timing shifts. Other attempts, such as ASAP [13] with human-assisted beat correction, and advancements in automatic beat tracking, such as Beat This! [14], aim to bring annotation to live-recorded datasets. However, accuracy limitations still present challenges.
@@ II-B1 Explicit Structural Features
The explicit use of structural features often relies on handcrafted feature engineering and external analysis tools. A common method in these models is a waterfall-like approach. Typically, the process begins by generating a lead sheet and then using the lead sheet as a condition for the subsequent generation tasks.
@@ II-B2 Implicit Structural Features
II-C Unconditional Generation for Full-Length Music
In the field of symbolic music generation, most studies adopt one or both of two primary paradigms: conditional and unconditional generation. For example, Music Transformer [7] employs both approaches. In a narrow sense, conditional generation refers to using a short segment of an existing music sequence as a prompt, which serves as the context for generating new tokens. The model then generates new tokens autoregressively based on this prompt, either by appending them to form a longer sequence or by extracting the generated continuation as the final output.
However, in such cases, the result is either not fully generated by the model if the prompt is included in the output, or, if only the continuation beyond the prompt is used, the generated segment lacks a clear and intentional beginning. This limitation becomes particularly problematic in full-length music generation, where human-composed works typically exhibit deliberate harmonic and melodic planning from a well-defined starting point. As highlighted in the work by Dai et al. [22], the section structure is not merely an arbitrary construction, but reflects significant interactions with harmony, melody, rhythm, and predictivity. Consequently, unconditional generation, in which the model produces the entire sequence from scratch without any initial prompt, is essential for achieving structurally complete music with a clear beginning and end.
It is worth noting that broader definitions of conditional generation may also include non-prompt-based conditioning or other control signals. However, such cases fall outside the scope of this study, which specifically aims to investigate the model’s ability to generate complete musical works under unconditional settings.
II-D SOTA Solutions with Long-Term Dependency on Performance Datasets
@@
When using a context length of , the standard Transformer incurs the following computational and memory costs for self-attention: FLOPs is , and memory usage is attention scores.
In contrast, Perceiver AR applies cross-attention from a much shorter latent sequence, significantly reducing the cost of subsequent self-attention. For example, if the latent length is , the self-attention within the latent space requires FLOPs of , and memory usage of attention scores.
This example shows that Perceiver AR can substantially reduce both computational cost and memory usage. Compared to the standard Transformer, the reduction is approximately 1,024 times in computation and 1,024 times in memory consumption.
However, Perceiver AR leverages the last tokens as the Query with a limited causal mask, and training with teacher-forcing on long sequences led to lower quality generation. Furthermore, we observed that relying solely on long, especially ultra-long, context resulted in repetitive segments in the latter part of the generated content.
Despite significant progress, existing research has yet to fully address the challenge of generating music with both long-term dependencies and expressive performance details, as most approaches either rely heavily on manual processes or face performance limitations due to restricted context window lengths.
Among the state-of-the-art models discussed above, Perceiver AR stands out as the one that employs the longest token context length in its experiments. It is also trained directly on performance data and does not require annotations such as beat, bar, or phrase. These characteristics make it uniquely well-suited as a baseline for our goal of learning to generate full-length music with structural completeness and expressive nuance from raw performance data.
@@ III-A2 Perceiver AR’s Causal Mask Issue
In Perceiver AR [4], the situation is different because the query token length is much smaller than the key and value lengths . Specifically, the causal mask only works on the final part of the context sequence, equivalent to the length of the query , but does not mask tokens that occur before that. This results in some tokens before the query length being visible during training, which is not an issue for generation, except that the segment from to is not properly learned by the model.
@@
Note that this issue primarily affects unconditional generation, where the model generates music without any specific prompt or primer, making it more sensitive to inconsistencies between training and generation contexts.
In conditional generation, where the model starts with an initial prompt or primer sequence, using a primer with a length close to the unmasked portion during training can help preserve generation quality and prevent degradation.
@@ III-B Ultra-Long Context in auto-regressive Generation
During generation, identical values are not produced at nearby steps to avoid training penalties. In fact, a similar context of neighboring tokens causes the generation process to produce identical or similar tokens at nearby steps, leading to a higher probability of generating repetitive short segments as the sequence grows longer.
In contrast, a short-range context attention mechanism that focuses only on the most recent tokens (e.g., ) changes the conditional probability distribution from:
This shorter context reduces the overlap between the context of and that of . As a result, the KL divergence between their conditional distributions increases:
This increase in divergence reduces the likelihood that , thereby lowering the token autocorrelation . Specifically, the term in Equation (9) becomes smaller on average due to decreased similarity between neighboring token predictions. Therefore, the introduction of short-range attention suppresses repetitive short segments and improves local diversity in the generated music.
@@ IV-A Improving the Model to Effectively Learn Ultra-Long Sequences
The upper part of Figure 1 demonstrates the original input sequence pre-processing of the baseline model. Specifically, the baseline model segments the input sequence using the maximum input length as the window size, leaving the beginning of the sequence uncovered by the causal mask. This results in the initial tokens not being predicted or supervised during teacher forcing, preventing the model from learning to generate accurate tokens at the beginning of the sequence due to their exclusion by the causal mask.
@@ V-A Effective Segmentation
This approach ensures that the model learns the sequence structure progressively, closely resembling the auto-regressive generation process, where tokens are predicted one at a time. As a result, the model is better aligned with the auto-regressive nature of music generation, reducing the peeking problem and improving generation quality. Thus, the model can smoothly learn the generation of sequences with an effective causal mask. The key advantages are summarized as follows.
@@ V-B Multi-Scale Cross-Attention Mechanism
First, we introduce the Attention mask without scale mask. In this case, all tokens are visible to the model, as defined in Equation (15):
@@ VI-A3 Hyperparameter Selection
The hidden dimension was set to 1,024 with 24 self-attention layers. Each attention layer had 16 heads, and the head dimension was 64. Adam optimization was used, with an initial learning rate set to 0.03125. Each generated sequence length was set to 8,192 tokens, resulting in approximately 2-10 minutes of music. The music generation in this setup was unconditional, meaning that no external conditions or prompts were used to guide the generation process.
In conditional generation, if the prompt length equals or exceeds the model’s maximum context window, the quality degradation typically caused by the limitations of the causal mask disappears. However, our goal is to generate full-length music with a natural beginning as part of a complete structure through unconditional generation. Therefore, using prompt-based settings that circumvent the limitations of causal masking does not align with the objectives of this study.
It is worth noting that the original Perceiver AR paper [4] briefly mentions the configuration used for long-context symbolic music generation as “1024 latents and 24 self-attention layers on input contexts of 32,768 tokens.” Beyond this, no further implementation details are disclosed. To ensure a fair comparison in this study, we adhere strictly to these disclosed parameters across both the baseline and improved models. Specifically, all models use a hidden dimension of 1024, 24 layers of latent self-attention, and 2 layers of cross-attention. The total number of parameters for the baseline and both improved models is matched at 361,621,520.
All models were trained on the Maestro dataset using a batch size of 1, a maximum sequence length of 32,768 tokens, a query length of 1,024, and the Adam optimizer with a learning rate following the Noam scheduler with 4,000 warmup steps. Each model was trained for 101 epochs, using cross-entropy loss for both training and validation. The training process exhibited clear validation loss stabilization and overall loss convergence before completion.
For sampling, we use the same decoding parameters for all models: top-p set to 0.75 and temperature set to 1.3. While reducing temperature can indeed help suppress erratic token generation and improve the musicality of the output, particularly in the baseline model, it comes at a cost. Lower temperature reduces diversity, leading to more repetitive and uniform outputs, a common issue in many current state-of-the-art music generation models. For this reason, we deliberately refrain from relying on artificially low temperatures to compensate for the baseline’s deficiencies.
Finally, 28 MIDI files generated by each model were rendered into audio using the Vintage Piano instrument from Logic Pro for subsequent evaluation.
@@ VI-A4 Evaluation Metrics
To calculate these additional segment-based metrics, each evaluation piece was divided into 64 segments of equal duration. All of these metrics are calculated based on the inter-set distribution similarity between the generated dataset and the ground truth dataset, including two values: the KL Divergence (KLD) and Overlap Area (OA). OA refers to the overlap between two probability density functions (PDFs), defined as the integral of the pointwise minimum of the two distributions. This value ranges from 0 (no overlap) to 1 (complete overlap), and it quantifies the similarity between two distributions. When the OA value is larger and the KLD value is smaller, it indicates that the above metric distributions of the generated dataset and the ground truth dataset are closely aligned, suggesting that the model has generated music more closely resembling human-composed and performed pieces.
@@ VI-B1 Experiment 1: Input Sequence Segmentation
Further Training of the Baseline Model under Different Hyperparameters
In the above experiment, we trained the baseline model using our current configuration (maximum input length of 32,768, query length of 1,024, and the conventional segmentation approach that randomly selects a starting point and crops a segment of max input length from the original sequence). We observed that the baseline model produced musically incoherent and chaotic outputs. To further investigate the behavior of the baseline model, we trained it under various maximum input length settings, specifically using 1,024, 2,048, 4,096, 8,196, and 16,384 tokens, in addition to the original 32,768-token setup. We then compared these six baseline models against our improved model that incorporates the proposed segmentation strategy.
By examining the validation loss curves of all models, we observed a consistent trend: the larger the gap between max input length and query length, the more difficult it became for the model’s validation loss to converge (see Figure 3). This result supports our earlier hypothesis that as the difference between the maximum input length and query length increases, the portion of tokens not covered by the causal mask and therefore not contributing to the loss during training also increases. This results in a higher ratio of untrained tokens, which hampers the model’s ability to generalize. Notably, under the extreme setting of a maximum input length of 1,024 and 2,048, the validation loss exhibited a decreasing trend, demonstrating that our baseline training procedure is itself valid and correctly implemented, and that the model’s degraded performance under other settings is not due to a flawed training methodology but rather to architectural and segmentation mismatches.
To further quantify this effect, we measured the proportion of tokens in the MAESTRO dataset [5] that could not be effectively utilized for training under different configurations. The total number of trainable tokens was calculated as the sum of the lengths of all token sequences in the dataset. For each max input length setting, we computed the number of unusable tokens per sequence as min(max input length, sequence length) - query length, representing the portion of tokens excluded from contributing to the model’s learning due to the causal attention constraint. The ratio between the excluded tokens and the total tokens was then used to estimate the inefficiency caused by the model design. As shown in Table II, the greater the difference between max input length and query length, the smaller the proportion of tokens that contribute to learning, corroborating our earlier claim that baseline models struggle to learn when the effective training signal is limited. These findings further support our proposed segmentation strategy as a necessary improvement for enabling better model generalization.
| Cell = Non-Contribution Tokens (Ineffective Token Rate) | ||||||
| Columns = Max Input Length | ||||||
| Query Length | 1,024 | 2,048 | 4,096 | 8,192 | 16,384 | 32,768 |
| 1,024 | 0 (0.00%) | 1,151,107 (4.45%) | 3,435,859 (13.28%) | 7,683,665 (29.71%) | 13,891,909 (53.71%) | 20,554,122 (79.47%) |
| 2,048 | 0 (0.00%) | 0 (0.00%) | 2,284,752 (8.83%) | 6,532,558 (25.26%) | 12,740,802 (49.26%) | 19,403,015 (75.02%) |
| 4,096 | 0 (0.00%) | 0 (0.00%) | 0 (0.00%) | 4,247,806 (16.42%) | 10,456,050 (40.43%) | 17,118,263 (66.19%) |
| 8,192 | 0 (0.00%) | 0 (0.00%) | 0 (0.00%) | 0 (0.00%) | 6,208,244 (24.00%) | 12,870,457 (49.76%) |
| 16,384 | 0 (0.00%) | 0 (0.00%) | 0 (0.00%) | 0 (0.00%) | 0 (0.00%) | 6,662,213 (25.76%) |
| 32,768 | 0 (0.00%) | 0 (0.00%) | 0 (0.00%) | 0 (0.00%) | 0 (0.00%) | 0 (0.00%) |
Note: Each cell shows the number of non-contributing tokens (excluded due to causal mask limits) and the corresponding ineffective token rate in parentheses. The total number of tokens in the Maestro dataset [5] is 25,864,128.
Given this analysis, the degraded generation quality of the baseline model under the (32,768, 1,024) setting becomes understandable. The original Perceiver AR paper [4] also found that increasing the context length on the Wikitext-103 dataset, especially to 4,096 and 8,192, led to degraded perplexity, which is consistent with the results of our experiments. To address this, we propose that segmentation-aware training is especially crucial when working with ultra-long contexts.